Make the subject of the formula S=ut+1/2 at2


1
Expert's answer
2019-11-12T12:29:29-0500

"S=ut+\frac{1}{2}at^2\\u=\frac{1}{t}(S-\frac{1}{2}at^2)\\a=\frac{2(S-ut)}{t^2}\\t=\frac{-u+\sqrt{u^2+2aS}}{a}\;or\;t=\frac{-u-\sqrt{u^2+2aS}}{a}"

Since "t\ge0", if a>0, only the first solution is valid, if a<0 both solutions are valid.